Ref Check for L^{A}TEX 2.09

This demo file is to describe usage of RefCheck package with L^{A}TEX 2.09. The package looks for useless labels, unlabelled equations and unused bibliography references.

To start checking one should just write

     \documentstyle[refcheck]{article}

at the beginning of document and compile the file twice. Then, comments on references could be found in LOG file.

Consider the set of equations

a = b  (1) eq.1
b = c  (2) ?\underline{eq.2}?
c = d,  (3) ???

written as

     \begin{eqnarray}
     &&a=b\label{eq.1}\\
     &&b=c\label{eq.2}\\
     &&c=d,
     \end{eqnarray}

Then the LOG file will contains the following messages:

 LaTeX Warning: Unused label `eq.2' on page 1 on input line 33.
 LaTeX Warning: Unlabelled equation (3) on page 1 on input line 35.

because the equation (1) is referenced here by (\ref{eq.1}), the label eq.2 of the second equation isn’t used and the last equation is numbered and unlabelled.

The labels identifiers on the marginal notes are stated by the command

     \showrefnames

These marks include some information on labels. Namely, the useless labels are underlined and bounded by ‘?’, the mark ‘???’ means that the equation is unlabelled. The marks are framed for labels which are used in text. This regime can be set off in any point of text by the command

     \norefnames

The package is working in the same way with bibliography. Consider the short list of books:

References

book1 [1] A book.

book2 [2] Another book.

?\underline{book3}? [3] Third book.

defined as follows

     \begin{thebibliography}{9}
     \bibitem{book1}A book.
     \bibitem{book2}Another book.
     \bibitem{book3}Third book.
     \end{thebibliography}

Further, we use the labels of [1, 2] with the command \cite{book1,book2} and do not use the label ‘book3’. Then, the LOG file will contain the following message

 LaTeX Warning: Unused bibitem `book3' on page 2 on input line 79.

Commands \showcitenames and \nocitenames for bibliography are analogous to the commands \showrefnames and \norefnames. They set on and off TEX to write to output labels of bibitems.

\underline{Note} the problem with implicitly cited labels. For example, if there is a set of equations which are labelled with a1,a2,...an and referred in

document by (\ref{a1})--(\ref{an}), then the labels a2,a3,...an-1

cannot be recognized to be ‘used’.